Your parents will retire in 21 years. They currently have $230,000 saved, and they think they will need $1,750,000 at retirement. What annual interest rate must they earn to reach their goal, assuming they don't save any additional funds? Round your answer to two decimal places.

  %
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,750,000=230,000*(1+r/100)^21

(1,750,000/230,000)^(1/21)=(1+r/100)

(1+r/100)=1.1015

r=1.1015-1

=10.15%(Approx).